OpenStudy (anonymous):

opens a savings account with 44.50. every week 22.25 is put into it. wants to buy a $178 item. how many weeks until you can buy it?

OpenStudy (kropot72):

178 = 44.5 + 22.25x where x is the number of weeks needed to save $178. You need to solve the equation to find x. Can you do that?

OpenStudy (anonymous):

i'll try:/

OpenStudy (anonymous):

i tried, cant do it

OpenStudy (anonymous):

Same time tommorow?:) i stay afterschool on weekdays soo @sriramkumar and @kropot72, thanks i think i got the right answer:p

OpenStudy (kropot72):

178 = 44.5 + 22.25x .................(1) The first step is to subtract 44.5 from both sides of equation (1) giving 178 - 44.5 = 44.5 - 44.5 + 22.25x ............(2) Equation (2) simplfies to 133.5 = 22.25x ..................(3) If you divide both sides of equation (3) by 22.25 you will get the value of x.

OpenStudy (anonymous):

the same one, I was intended to tell @kropot72 smart :D